OpenStudy (andertj21):

What is the solution set of 4x 2 - 36 = 0? {-3} {3} {-3, 3}

OpenStudy (anonymous):

is that 4x^2

OpenStudy (andertj21):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

LAs tone

OpenStudy (anonymous):

Last one

OpenStudy (anonymous):

4x^2-36=0 4x^2=36 2x=6 x=3

OpenStudy (anonymous):

(-3,3)

OpenStudy (andertj21):

thanks!

OpenStudy (anonymous):

you do know its the second one -.- i just did it

OpenStudy (anonymous):

Plug in -3 and then 3

OpenStudy (anonymous):

@Official_Hotboy

OpenStudy (anonymous):

you cant square root a negetive

OpenStudy (anonymous):

You can square it though

OpenStudy (anonymous):

-12^2 = 144 144-36 dose not = 0

OpenStudy (anonymous):

cat got ur tonge

OpenStudy (anonymous):

Order of Operations Exponents come first not multiplication

OpenStudy (anonymous):

the equation is 4x^2-36=0 not 4x(2)-26=0

OpenStudy (anonymous):

36*

OpenStudy (anonymous):

I know exponents=power

OpenStudy (anonymous):

.... im wasting my time here i worked it out already so just refrence tht

OpenStudy (anonymous):

Plug it in a calculator it might help you out a lot